Month: October 2005

New News Cluster.

Fri Oct 28 10:22:20 PDT 2005 — New News Cluster. Our new news cluster will be put into service early next week. It is a massive upgrade from the existing, built on a total of six AMD Opteron servers with nearly 30 Terabyte of local spool storage on over 90 SCSI and SATA disks. In addition to the local spools, the news reader servers have full access to a commercial news provider to ensure full completion. For more details on the new news cluster’s hardware and configuration please see news://

An brief history of our local news servers can be found in the following MOTD entries. Every disk in the new spool servers is larger than the complete capacity of the current news server.

Shortly after the new news cluster has been promoted to handle, access to will be discontinued. We also expect to discontinue access to in the near future but do not yet have a firm date set.

While we have tried to keep article numbering matched on the new cluster you may need to unsubscribe and resubscribe to get the correct list of articles from the server. This should not be a problem if you only read sonic.*, ba.* or sonoma.* where the full archives have been migrated from the old server.

Special thanks is due to Joe Greco for all his help with the turn up of the new cluster. -Kelsey and Nathan

DSL outage update.

Wed Oct 26 12:49:37 PDT 2005 — DSL outage update. Service has been restored to most subscribers in the Santa Rosa area. Sync-no-surf persists in Mountain View. ASI reports similar outages at service centers throughout the greater Bay Area. We have no ETR at this time. -John F and Tech Support

Partial DSL outage.

Wed Oct 26 11:13:38 PDT 2005 — Partial DSL outage. We are receiving reports of sync-no-surf outages affecting parts of Santa Rosa and Mountain View. Not all customers are affected. We are in contact with SBC and are working on a resolution. -Support

System maintenance.

Tue Oct 18 10:55:17 PDT 2005 — System maintenance. At 12:01 AM Wednesday, Oct 19, we will be rebooting the systems that serve as the head-end of our Hotspot wireless network. DSL customers hosting Hotspots will experience approximately 5 minutes of downtime.

News server issues.

Tue Oct 18 09:21:49 PDT 2005 — News server issues. Last night news feeds to our local news server,, ran into some issues and stopped. We are currently back-feeding articles to the server. Hopefully little, if any news will be lost. -Kelsey

Office phone system restored.

Sun Oct 16 19:00:32 PDT 2005 — Office phone system restored. Telephone service to the offices (including Technical Support) was restored at approximately 6:56pm. Thank you for your patience. – John F

Office phone system outage.

Sun Oct 16 17:50:12 PDT 2005 — Office phone system outage. Today at approximately 5:08pm, the office phone system went offline. This only affects our office telephone system, and should have no impact on customer connectivity. We will update this message with more information as it becomes available. – John F

SMS Reboot.

Wed Oct 12 12:06:30 PDT 2005 — SMS Reboot. The router that terminates DSL connections in LATA9 (Stockton area) has just rebooted. We are investigating the cause of the reboot. All customers that were affected by this reboot are back online.

Core Switch Maintenance.

Fri Oct 7 08:19:35 PDT 2005 — Core Switch Maintenance. On Saturday 10/8 at 12:01AM we will be replacing a switching module in one of our core switches at our Santa Rosa datacenter. This operation may result in a brief loss of connectivity to certain destinations for a period of approximately 5 minutes while we perform the necessary hardware swap. – Network Operations

Correction: Problems on numbers ending in 5.

Wed Oct 5 19:10:13 PDT 2005 — Correction: Problems on numbers ending in 5. The access numbers experiencing problems all end in 5. Numbers ending in 3 are not experiencing problems. — Russ

Update Wed Oct 5 21:26:15 PDT 2005 — The problems with the dial group has been resolved. -Russ